### v1

# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

Abstract
Off-policy learning serves as the primary framework for learning optimal policies from logged interactions collected under a static behavior policy. In this work, we investigate the more practical and flexible setting of adaptive off-policy learning, where policies are iteratively refined and re-deployed to collect higher-quality data. Building on the success of PAC-Bayesian learning with Logarithmic Smoothing (LS) in static settings, we extend this framework to the adaptive scenario using tools from online PAC-Bayesian theory. Furthermore, we demonstrate that a principled adjustment to the LS estimator naturally accommodates multiple rounds of deployment and yields faster convergence rates under mild conditions. Our method matches the performance of leading offline approaches in static settings, and significantly outperforms them when intermediate policy deployments are allowed. Empirical evaluations across diverse scenarios highlight both the advantages of adaptive data collection and the strength of the PAC-Bayesian formulation.
1 Introduction
Off-policy learning has become a ubiquitous task for practitioners in a wide range of applications such as recommender systems (Sakhi_Rohde_Gilotte_2023, ), ad placement (bottou2013counterfactual, ), and personalized medicine (kallus2018policy, ). This task involves learning optimal policies from offline data summarizing a behavior policy’s past interactions. These interactions, typically captured as tuples of contextual features, action played, and reward received, hold valuable insight into the underlying dynamics of the environment. Each tuple represents a single round of interaction, where the agent observes a context (including relevant features), takes an action according to its current, behavior policy, and receives a reward that depends on both the observed context and the taken action.
The off-policy learning literature focused on developing efficient approaches to learn from these logged interactions bottou2013counterfactual ; swaminathan2015batch ; london2019bayesian ; sakhi2023pac . At the core of these methods the development of tailored off-policy estimators with improved concentration properties (bottou2013counterfactual, ; aouali23a, ; metelli2021subgaussian, ; sakhi2024logarithmic, ), combined with tight generalization bounds allowing efficient learning (swaminathan2015batch, ; sakhi2023pac, ; sakhi2024logarithmic, ). The methodologies employed derive from the pessimism principle jin2021pessimism , which has been demonstrated to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens london2019bayesian proved efficient, improving the seminal work of (swaminathan2015batch, ). This direction was further explored by sakhi2023pac ; aouali23a ; pac_bayes_unified ; gabbianelli2023importance ; sakhi2024logarithmic , reducing assumptions and improving the pessimistic principle, i.e. controlling the worst-case scenario, with latest developments (sakhi2024logarithmic, ) resulting in state of the art pessimistic approaches.
The success of these pessimistic learning approaches fundamentally depend on the quality of the logged data (jin2021pessimism, ; gabbianelli2023importance, ; sakhi2024logarithmic, ), thus the quality of the behavior policy and its ability to play meaningful actions. In particular, it was demonstrated that off-policy learning approaches converge to the optimal policy with enough samples, with faster convergence when the behavior policy explores well the optimal actions (sakhi2024logarithmic, ). To achieve this convergence, a naive strategy consists in collecting more data with the same logging system. However, this approach is limited as it does not consider the benefits of learning from previous interactions. Then, a better adaptive strategy would use the already collected data to refine the design of the behavior policy, adapting better to the underlying dynamics of the problem, and collecting interaction data of higher quality. More precisely, the adaptive strategy performs a sequence of data collection experiments by determining at each round a new behavior policy using interaction data collected during previous experiments. The obtained policy is then deployed for the next round to collect better quality samples.
This adaptive (or sequential) off-policy learning approach recently sparked interest (adaptive, ; scrm, ). However, existing algorithms are not in line with common practices, either requiring strong assumptions such as uniform coverage adaptive ; scrm , or involving unrealistic data collection strategies and cannot accommodate online updates scrm . In the conventional off-policy learning setting, a prominent route exploits PAC-Bayes learning guedj2019primer ; hellstrom2023generalization ; alquier2024user to reach both theory-driven learning algorithm and convergence guarantees sakhi2024logarithmic . Given the existence of online PAC-Bayes methods via supermartingales haddouche2022online ; haddouche2023pac ; chugg2023unified allowing a continuous data stream, a natural question then arise: can we reach PAC-Bayesian adaptive off-policy algorithms?
Intuitively, an adaptive PAC-Bayesian strategy would yield in better behavior policies at each deployment than sakhi2024logarithmic , increasing the quality of the logged samples, which in turn result in a faster convergence towards the optimal policy. In this work, we formalize this intuition and develop the necessary tools to study this strategy and obtain such learning guarantees.
Contributions. By leveraging the online PAC-Bayesian toolbox, we extend in Sec. 3 the offline pessimistic logarithmic smoothing PAC-Bayesian strategy sakhi2024logarithmic to the adaptive scenario adaptive , allowing the approach to handle sequential interaction data, collected by progressively adaptive behavior policies. This helps us define a first algorithm that already exhibits improved performance compared to the offline approach. These improvements do not require sophisticated data collection strategies nor strong assumptions. Additionally, we identify a core limitation of the logarithmic smoothing estimator and propose in Sec. 4 an adjustment to address it. This correction allows for faster convergence rates to the optimal policy under reasonable and interpretable assumptions. Our proposed approach matches the guarantees of the offline setting and improves upon them when we have the possibility to deploy multiple policies, completely generalizing the state-of-the-art offline PAC-Bayesian approach and allowing for adaptive PAC-Bayesian off-policy learning. We also conduct in Sec. 5 numerical experiments to demonstrate the effectiveness of our approach and quantify how it improves upon the pure offline setting as well as recent adaptive strategies.
Outline. Sec. 2 introduces the adaptive framework and provides the necessary background. In Sec. 3, we provide a first extension to the logarithmic smoothing (LS) PAC-Bayesian off-policy learning approach to the adaptive setting, study its benefits and limitation. In Sec. 4, we provide a simple adjustment to the LS estimator, achieving an accelerated convergence rate under mild conditions. We also study its behavior under different data collection strategies and demonstrate its effectiveness for all scenarios. Empirical validation is conducted in Sec. 5 to highlight the favorable performance of the our proposed method, and Sec. 6 provides concluding remarks.
2 Framework
Adaptive Contextual bandits.
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the context space, which is a compact subset of $\mathbb{R}^{d}$, and let $\mathcal{A}=[K]$ be a finite action set. An agent’s actions are guided by a stochastic policy $\pi:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{A})$ within a policy space $\Pi$. Given a context $x\in\mathcal{X},\pi(\cdot\mid x)$ is a probability distribution over the action set $\mathcal{A};\pi(a\mid x)$ is the probability that the agent selects action $a$ in context $x$. In a general adaptive framework, we consider that the agent interacts sequentially with a data stream: at time $k\geq 0$, the agent deploys a new behavior policy $\pi_{k}$ and a new batch of contexts $\{x_{j}\mid j\in I_{k}\}$ observed by the agent is revealed, where $|I_{k}|=n_{k}$ and all $(I_{k})_{k\geq 0}$ are disjoint. These contexts are drawn i.i.d. from a fixed, unknown distribution $\nu$. For each $j\in I_{k}$, the agent selects for context $x_{j}$ an action $a_{j}\sim\pi_{k}\left(\cdot\mid x_{j}\right)$, where $\pi_{k}$ is the current behavior policy of the agent at time $k$, and receives a stochastic cost $c_{j}\in[-1,0]$ that depends on the observed context $x_{j}$ and the taken action $a_{j}$. This cost $c_{j}$ is sampled from an unknown cost distribution $p\left(\cdot|x_{j},a_{j}\right)$ and we denote by $c(a,x)$ its expectation under $p\left(\cdot|x_{j},a_{j}\right)$. After $|I_{k}|=n_{k}$ interactions, the agent adaptively updates its behavior to $\pi_{k+1}$ w.r.t. previous behavior policies $(\pi_{j})_{j=0\cdots k}$ and is allowed to use all available data $\{x_{i},a_{i},c_{i}\}_{i=1\cdots N_{k}}$, where $N_{k}:=\sum_{i=0}^{k}n_{i}$. To ease notations, we denote by $P(\pi):=\nu\times\pi\times p$ the distribution of $(x,a,c)$ tuples generated by policy $\pi$. The performance of any policy $\pi$ is assessed by its risk, which is defined as the expectation under the context distribution $\nu$, and the played actions $a$ by $\pi$ of the expected costs $c(x,a)$, as in:
$R(\pi)=\mathbb{E}_{(x,a,c)\sim P(\pi)}[c]=\mathbb{E}_{x\sim\nu,a\sim\pi(\cdot\mid x)}[c(x,a)]\,.$
In this work, our goal is to come up with adaptive learning algorithms that guarantee the convergence to the optimal policy $\pi_{\star}$, defined as the policy $\pi$ minimizing the risk as: $\pi_{\star}=\operatorname*{arg\,min}_{\pi\in\Pi}R(\pi)\,.$
PAC-Bayes learning. To learn the optimal policy $\pi_{\star}$, we employ the pessimism principle (jin2021pessimism, ), shown to be optimal for off-policy learning jin2021pessimism . Specifically, we aim at minimizing an empirical, tight upper bound on the risk $R(\pi)$ over policies $\pi\in\Pi$. To do so, we leverage PAC-Bayes learning (alquier2024user, ) which controls the generalization guarantee of a posterior distribution $Q$ over a predictor space $\Theta$ (namely $Q\in\mathcal{P}(\Theta)$). To make this consistent with controlling the risk of a policy $\pi$, we exploit a specific parametrization of policies $\pi\in\Pi$. For any $\theta\in\Theta$, we define the deterministic policy: $\forall x,d_{\theta}(x):=\mathds{1}\{a=f_{\theta}(x)\}$, where $f_{\theta}:\mathcal{X}\rightarrow\mathcal{A}$, a parametrized predictor. Then, each distribution $Q\in\mathcal{P}(\Theta)$ defines a policy $\pi_{Q}$ by setting the probability of action $a$ given context $x$ as the probability that a random predictor $f_{\theta}\sim Q$ maps $x$ to action $a$, that is:
$\forall(x,a)\in\mathcal{X}\times\mathcal{A}\,,\quad\pi_{Q}(a|x)=\mathbb{E}_{\theta\sim Q}\left[\mathds{1}\{a=f_{\theta}(x)\}\right]\,.$
This expression is not restrictive as any policy can be represented in this form (sakhi2023pac, , Theorem 3.1). With this characterization, the space of policies considered $\Pi$ can be entirely described by a predictor space $\Theta$ and a set of distributions over it $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$ sakhi2023pac . With this duality, we can obtain with the PAC-Bayesian toolbox, by fixing a reference distribution $P$, and setting $\delta\in(0,1]$ a bound of the following general form, holding with probability $1-\delta$, simultaneously for all $Q$:
$\forall Q\in\mathcal{C}(\Theta)\,,R(\pi_{Q})\leq\hat{R}(\pi_{Q})+\mathcal{O}\left(\operatorname{KL}(Q||P)\right)\,,$
upper bounding the true risk of $\pi_{Q}$ by a a risk estimator $\hat{R}(\pi_{Q})$ and $\operatorname{KL}(Q||P):=\mathbb{E}_{Q}\left[\log\left(\frac{dQ}{dP}\right)\right]$ is the Kullback-Leibler divergence between $P$ and $Q$. The tightness of the bound fundamentally depends on the concentration properties of the chosen estimator $\hat{R}$ sakhi2024logarithmic .
Adaptive algorithms. We define the full (countable) dataset $S:=\cup_{k\in\mathbb{N}}S_{k}$, where $S_{k}:=\left\{(x_{j},a_{j},c_{j})\mid j\in I_{k}\right\}$ is the set of logged data at time $k$. We also define $\mathcal{F}_{k}$ a filtration adapted to $\left(\cup_{j=0}^{k}S_{j}\right)_{j\in 0\cdots k}$ ($\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra). This framework generalizes the pure static case, on which is based sakhi2024logarithmic , as $S_{0}$ summarizes the interactions of a single data batch with the starting policy $\pi_{0}$. We denote as adaptive algorithm, any algorithm Alg such that its output $\pi_{k+1}$ at time $k$ is $\mathcal{F}_{k}$-measurable. In other words, if we denote all the collected interactions at time $k$ as $S_{0:k}:=\cup_{\ell=0}^{k}S_{\ell}$, , we can express $\pi_{k+1}=\texttt{Alg}(S_{0:k})$. We assume that for each $k$, $(x_{j})_{j\in I_{k}}\sim\nu^{\otimes n_{k}}$ is i.i.d. and that, conditionally to $\mathcal{F}_{k-1}$, $S_{k}$ is constituted of mutually independent random variables.
Regularized IPS. Given the successive arrival of data batches $(S_{k})_{k\geq 0}$ and for all $j\geq 0$, $\pi_{j+1}=\texttt{Alg}(S_{0:j})$ is given by an adaptive algorithm, we can leverage the collected data at any time $k\geq 0$ to define an Inverse Propensity Scoring (IPS) estimator horvitz1952generalization , for the risk of any policy $\pi$: $\hat{R}_{0:k}^{\textsc{IPS}}(\pi)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}c_{i}\,,$ which recovers the static case when $k=0$. Even with the potential dependencies between the behavior policies $\pi_{j}$, this estimator is unbiased under the common support assumption mcbook . However, as in the static case, this estimator has poor concentration properties metelli2021subgaussian and regularized variants, trading some bias for lower variance are preferred bottou2013counterfactual ; optimistic_shrinkage ; aouali23a ; sakhi2024logarithmic . The majority of these regularized estimators can be defined by specific choices of a regularizer $h:[0,1]^{2}\times[-1,0]\rightarrow\mathbb{R}^{-}$ in the following general expression:
$\hat{R}_{0:k}^{h}(\pi)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}h(\pi(a_{i}|x_{i}),\pi_{j}(a_{i}|x_{i}),c_{i})\,,$
with $h$ respecting $\forall(p,q,c)\in[0,1]^{2}\times[-1,0]\,,\quad pc/q\leq h(p,q,c)\leq 0$. $h(p,q,c)=pc/q$ recovers the IPS estimator while other choices recover known, well behaved estimators such as clipped IPS bottou2013counterfactual , harmonic metelli2021subgaussian and Implicit-Exploration gabbianelli2023importance to name a few. When $k=0$, we denote $\hat{R}_{0:0}^{h}$ as $\hat{R}_{0}^{h}$.
3 Warm-up on Pessimistic Logarithmic Smoothing
3.1 Batch learning algorithms with theoretical guarantees
Pessimistic Logarithmic Smoothing recently appeared as a fruitful theoretical approach to obtain both concrete learning algorithms and generalization guarantees guarantees for contextual bandits. Those coupled benefits naturally stems from the use of PAC-Bayesian theory alquier2024user which provide empirical generalization bounds whom the optimization yield original learning algorithms with theoretical guarantees. In sakhi2024logarithmic , pessimistic PAC-Bayes bounds have been developed for contextual bandits, providing a novel learning algorithm with optimality guarantee. We first recall the definition of Logarithmic smoothing as well as its associated regularized IPS estimator. For any $\lambda>0,k\geq 0$,
$h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q}\right)\quad\text{and}\quad\hat{R}_{0:k}^{\lambda\mathrm{-LS}}:=\hat{R}_{0:k}^{h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}}.$
Using these notions, sakhi2024logarithmic obtained results for static off-policy learning i.e. data are gathered in a single batch $S_{0}$ of size $n_{0}$. We recall them below.
Proposition 3.1 (Propositions 10 and 11 of sakhi2024logarithmic, ).
Given a prior $P\in\mathcal{P}(\Theta),\delta\in(0,1]$ and $\lambda>0$, the following holds with probability at least $1-\delta$ over $S_{0}$:
$\forall Q\in\mathcal{P}(\Theta),\quad R\left(\pi_{Q}\right)\leq\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)+\log\frac{1}{\delta}}{\lambda n_{0}}.$
Furthermore, for any $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$, denoting by $Q_{n_{0}}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)}{\lambda n_{0}}$ and $Q^{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$, we have:
00 0 $\leq R(\pi_{Q_{n_{0}}})-R(\pi_{Q^{\star}})\leq\lambda\mathcal{S}(\pi_{Q^{\star}})+\frac{2\operatorname{KL}(Q^{\star}\|P)+\log\frac{2}{\delta}}{\lambda n_{0}},$
where $\mathcal{S}(\pi):=\mathbb{E}_{(x,a,c)\sim P(\pi_{0})}\left[\frac{\pi(a\mid x)c^{2}}{\pi_{0}^{2}(a\mid x)}\right]$ is a pseudo-variance term.
Those results provide two distinct hindsights: (i), the learning procedure yielding $Q_{n_{0}}$ (and exploited in practice in sakhi2024logarithmic ) minimises a generalisation bound and (ii), once $\mathcal{S}(\pi_{Q^{\star}})$ finite, taking $\lambda=1/\sqrt{n_{0}}$ gives a convergence rate of $1/\sqrt{n_{0}}$ to the optimal value of the risk. However, Proposition 3.1 and its associated learning algorithm suffers from theoretical and practical limitations.
  1. 1.
    A batch procedure cannot handle a progressive arrival of data $(S_{k})_{k\geq 0}$. In particular, Proposition 3.1 does not allow to re-train our policy w.r.t all past data when a new data batch arrives.
  2. 2.
    The pseudo-variance term of Proposition 3.1 does not go to zero, even in the oracle case where $\pi_{0}=\pi_{Q^{\star}}$, whatever the optimal policy $\pi_{Q^{\star}}$ is. This exhibits a fundamental limitation of this batch procedure as the slow convergence rate of $1/\sqrt{n_{0}}$ is inevitable.
We provide below alternatives for those two fundamental issues. To answer 1. we show in Sec. 3.2 that leveraging online PAC-Bayes and supermartingale techniques haddouche2022online ; haddouche2023pac it is possible to obtain an adaptive version of Proposition 3.1, allowing both the arrival of new data on-the-fly and re-using former ones. Then, re-using online PAC-Bayes techniques, we answer 2. in Sec. 4 by providing a novel adaptive learning algorithm with accelerated convergence rate.
3.2 Towards a first adaptive pessimistic learning algorithm
In this part, we assume, for pedagogical concerns, that all datasets $S_{k}$ have the same size $n_{k}=m$. Still exploiting Logarithmic Smoothing, we provide an adaptive counterpart of Proposition 3.1 by involving online PAC-Bayes learning techniques haddouche2022online ; haddouche2023pac . From a technical perpective, we allow the progressive arrival of data batches $(S_{k})_{k\geq 0}$ by controlling the expansion of a supermartingale through Ville’s inequality doob1939jean instead of an exponential moment via Markov’s inequality. Our result is stated as follows:
{restatable}
theoremextensionsakhi For any adaptive algorithm Alg, any $m>0$ and data independent prior $P$. For any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0,$ denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$R(\pi_{Q})\leq\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda(k+1)m}.$
The proof lies in Appendix D. Note that taking $k=0$ and $m=n_{0}$ retrieves exactly the batch case of Proposition 3.1. On the contrary taking $k=n$ and $m=1$ provides a purely online procedure. Our theorem provides an interpolation between batch and online scenarios, allowing sub-batches of data to arrive at each time step and optimize on-the-fly.
Notice that the optimization problem $\operatorname{argmin}_{Q\in\mathcal{P}(\Theta)}\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q})+\frac{\operatorname{KL}(Q,P)}{\lambda(k+1)m}$ admits a closed form solution: the Gibbs posterior $\mathrm{Q}_{\lambda}$ such that $dQ_{\lambda}(\theta)\propto\exp\left(-\lambda(k+1)m\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(d_{\theta})\right)$ (catoni2007pac, , Theorem 1.2.6). However, Gibbs posteriors often requires costly, advanced Monte Carlo methods mcbook ; smc_book to be computed. A common practice to approximate the Gibbs posterior is to restrict the optimization problem on a subclass $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$ on which the optimisation problem is tractable, a widely used class in PAC-Bayes learning is the class of Gaussian distributions dziugaite2017computing ; perez2021tighter ; haddouche2022online .
Finally, a key remark is that, given a fixed batch size $m>0$ and parameter $\lambda>0$, Sec. 3.2 holds for a countable number $k$ of sub-batches simultaneously. A fundamental difference with Proposition 3.1 is that we allow the empirical risk $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ to evolve with $k$, while the batch approach only tackles $S_{0}$. This is crucial as $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ that evolve adaptively to the intermediate policies $(\pi_{j})_{j=0\cdots k}$ while $\hat{R}_{0}^{\lambda-\mathrm{LS}}$ only involves the initial policy $\pi_{0}$, which does improve during the learning process. Thus, optimising Sec. 3.2 at each time step yields an adaptive pessimistic learning algorithm given in Algorithm 1.
1 Input: Behavior policy $\pi_{0}$, a policy class $\mathcal{C}(\Theta)$, $\lambda>0$.
2 Initialize $P$ such that $\pi_{P}=\pi_{0}$ and $\mathcal{S}_{-1}=\{\}$.
3 for $k\geq 0$ do
4       Collect data $S_{k}$ of size $|I_{k}|=m$ with $\pi_{k}$ and define global collection $S_{0:k}=S_{0:k-1}\cup S_{k}$.
5       Solve the optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\left\{\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda km}\right\}$ (1)
Set $\pi_{k+1}=\pi_{\hat{Q}_{k+1}}$.
Algorithm 1 Adaptive Policy Learning via Logarithmic Smoothing
Note that at time $k$, we instantiate $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ with $(\pi_{j})_{j=0\cdots k}$ the outputs of Algorithm 1 up to time $k-1$. Similarly to the batch case, Algorithm 1 also enjoys a convergence guarantee. {restatable}propositionslowconvergence For any fixed subbatch size $m>0$, any data independent prior $P$, any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$, for any $k\geq 0$, denoting by $\pi_{k+1}$ the output of Algorithm 1 at time $k$ and $\pi_{\star}=\pi_{Q_{\star}}$ with $Q_{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$:
$R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}\left(\pi_{\star}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
where $\mathcal{S}_{j}(\pi):=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\pi(a\mid x)c^{2}}{\pi_{j}^{2}(a\mid x)}\right]$. Algorithm 1 is proven in Appendix D, and taking $\lambda=\frac{1}{\sqrt{(k+1)m}}$ yields an overall convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{km}}\right)$, matching the rate of Proposition 3.1 in the batch case (i.e $k=0,m=n_{0}$) and yielding an original rate of $\mathcal{O}(1/\sqrt{n})$ for an online setting ($k=n,m=1$). Algorithm 1 also works for any intermediate regime between batch and online settings, showing the strength of the adaptive approach. Looking at Algorithm 1, we observe that the upper bound depends on the average $\frac{1}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}(\pi_{\star})$, which is expected to be smaller than $\mathcal{S}_{0}(\pi_{\star})$, as intermediate policies $\pi_{j}$ should be progressively closer to $\pi_{\star}$. However, this algorithm still suffers from the same limitations than the static case here as the pseudo-variance terms $\mathcal{S}_{j}$ do not decrease to zero, even in the perfect situation where for all $j$, $\pi_{j}=\pi_{\star}$.
In this part, we showed that online PAC-Bayes techniques successfully extend offline algorithms to adaptive ones, improving dependencies while enjoying the same convergence rates. A key question is then: is it possible to find an adaptive algorithm with an accelerated convergence rate?
4 Adaptive Logarithmic Smoothing with accelerated convergence.
While Sec. 3.2 provide a natural adaptive extension of sakhi2024logarithmic , Algorithm 1 still suffers (Algorithm 1) from a slow convergence rate of $\mathcal{O}(1/\sqrt{km})$ due to the nature of the pseudo-variance term. Is it then possible to avoid such a term? Yes, by adjusting $h^{\lambda-\mathrm{Lin}}$ introduced in Sec. 3: we define $h^{\text{adj}}_{\lambda}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q(1+\lambda c)}\right)$ and denote it as the adjusted linear LS regularizer. We then define the adjusted risk for any adaptive algorithm Alg and $\lambda\in(0,1),k\geq 0$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi):=R_{0,k}^{h^{\text{adj}}_{\lambda}}(\pi)=-\frac{1}{N_{k}}\sum_{j=1}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\lambda}\log\left(1-\frac{\lambda c_{i}}{\pi_{j}(a_{i}|x_{i})(1+\lambda c_{i})}\right),$ (2)
where again, $\pi_{k+1}$ is the output of Alg at time $k$. Using similar tools than Sec. 3.2, we are able to reach the following theoretical guarantee for our novel regularizer $h^{\text{adj}}_{\lambda}$
Theorem 4.1.
For any adaptive algorithm Alg, any sequence $(n_{k})_{k\geq 0}$ and data independent prior $P$. For any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$ $\leq\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}},$
with $\hat{C}_{0,k}(\lambda)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}\frac{1}{\lambda}\log(\frac{1}{1+\lambda c_{i}})$.
Thm. 4.1 is obtained in Appendix E, using the same tools than Sec. 3.2 and thus enjoys the same general remarks. Note that this result holds for varying batch sizes $(n_{k})_{k\geq 0}$, allowing to encompass real-life situations where there are discrepancies in the arrival of new data. In particular, it is the theoretical backbone behind the main algorithm of this work: an adaptive and adjusted learning algorithm for contextual bandits, stated in Algorithm 2.
1 Input: Behavior policy $\pi_{0}$, a policy class $\mathcal{C}(\Theta)$, $\lambda\in(0,1)$.
2 Initialize $P$ such that $\pi_{P}=\pi_{0}$ and $\mathcal{S}_{-1}=\{\}$.
3 for $k\geq 0$: do
4       Collect data $S_{k}$ of size $|I_{k}|=n_{k}$ with $\pi_{k}$.
5       Define global collection $S_{0:k}=S_{0:k-1}\cup S_{k}$, and $N_{k}:=|S_{0:k}|$.
6       Solve the following optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}}\left\{\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda N_{k}}\right\}$ (3)
Set $\pi_{k+1}=\pi_{\hat{Q}_{k+1}}$.
Algorithm 2 Adaptive Policy Learning via Adjusted Logarithmic Smoothing
Note that at time $k$, we instantiate $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})$ with $(\pi_{j})_{j=0\cdots k}$ the previous outputs of Algorithm 2.
Note that in Thm. 4.1, contrary to Sec. 3.2, we do not control directly $R(\pi_{Q})$ but the difference $R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$. This means that running Algorithm 2 at time $k$ using $\cup_{j=0}^{k}S_{j}$ tends to ensure a better generalisation ability than what we obtained (on a weighted average) by running Algorithm 2 on $\cup_{j=0}^{k-1}S_{j}$. This phenomenon gives an intuition about why Algorithm 2 converge quickly to the optimum. We prove this intuition in what follows.
Towards an accelerated convergence rate for Algorithm 2. Recall that $\Pi$ is the considered space of policies on which the optimization of the risk is conducted. If $\Pi$ is the space of all possible policies, the optimal policy $\pi_{\star}$ can be characterized. For any context $x$, let $a_{\star}(x)=\operatorname*{arg\,min}_{a^{\prime}\in\mathcal{A}}c(a^{\prime},x)$ with ties broken arbitrarily. We obtain the optimal policy $\pi_{\star}$ as the one that chooses deterministically the optimal action $a_{\star}(x)$ for any context $x$, giving:
$\forall(x,a)\in\mathcal{X}\times\mathcal{A}\,,\quad\pi_{\star}(a|x)=\mathds{1}\left[a=a_{\star}(x)\right].$
In the following, we assume that our class of parametrized predictors $\{f_{\theta}\mid\theta\in\Theta\}$ is rich enough to represent $\pi_{\star}$. Then, there exists a $Q^{\star}\in\mathcal{C}(\Theta)$ such that $\pi_{\star}=\pi_{Q^{\star}}$. To prove the fast convergence of Algorithm 2 to $\pi_{\star}$, we state a first proposition.
Proposition 4.2.
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$0\leq R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{1-\lambda}\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}},$
with $L(\pi,\pi_{j})=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\left(\pi(a|x)\left(\left(\frac{1}{\pi_{j}(a|x)-1}\right)^{2}-1\right)+1\right)c^{2}\right]\,.$
Proposition 4.2 is proven in Appendix E and highlights the interest of $h^{\text{adj}}_{\lambda}$: we avoid the pseudo variance terms $\mathcal{S}_{j}(\pi_{\star})$ of Algorithm 1 to get $L\left(\pi_{\star},\pi_{j}\right)$. Contrary to $S_{j}(\pi_{\star})$, we have that, for deterministic policies $\pi_{\star}$, $L(\pi_{\star},\pi_{j})\rightarrow 0$ if $\pi_{j}\rightarrow\pi_{\star}$. Such desirable behavior leaves room for accelerated convergence rate. To investigate this route, we target the convergence towards the optimal policy $\pi_{\star}$ over the whole space and study the behavior of $L$.
To prove the acceleration of our method, we need to define additional quantities. We define the action $a$ sub-optimality given a context $x$ by: $\Delta_{a,x}:=c(a,x)-c(a_{\star}(x),x)\geq 0$. Then, for any $u\in[0,1)$, we define $\Delta_{u}$ such that:
$P_{x}\left(\min_{a\neq a^{\star}(x)}\Delta_{a,x}\geq\Delta_{u}\right)=1-u.$
For a fixed $u$, $\Delta_{u}$ quantifies how separable $a^{\star}(x)$ is from the rest of actions. In other words, the magnitude of $\Delta_{u}$ quantifies the difficulty of the problem. $\Delta_{u}=0$ for all $u\in[0,1)$ means that the optimal action $a^{\star}(x)$ is not distinguishable from the second best action, almost surely. Oppositely, $\Delta_{u}=1$ for a specific $u$ means that with probability $1-u$, the costs are deterministic and the optimal action $a^{\star}(x)$ is the only one to have a non null cost (of $-1$). We want to avoid the first case, as the existence of a strictly positive $\Delta_{u}>0$ will be important next. Finally, we introduce the minimal plausibility of the optimal actions under $\pi_{k}$:
$C^{\star}_{k}=\inf_{x}\pi_{k}(a^{\star}(x)|x)\,.$
Lemma 4.3.
Let $\pi_{k}\in\Pi$, assume there exists $u\in[0,1)$ such that $\Delta_{u}>0$ and $C^{\star}_{k}>0$, then:
$L(\pi_{\star},\pi_{k})\leq\gamma_{k}(R(\pi_{k})-R(\pi_{\star})),$
with $\gamma_{k}=1+\left(\frac{1}{4}+\frac{1}{C_{k}^{\star}}\right)\frac{1}{\Delta_{u}(1-u)}=\mathcal{O}\left(\frac{1}{(1-u)\Delta_{u}C^{\star}_{k}}\right).$
Proof is in Appendix E. The existence of a $\Delta_{u}>0$ is similar to margin conditions used in multi-arm contextual bandits fast_rate_cb ; adaptive , and $C^{\star}_{k}>0$ is a condition on the coverage of the optimal actions by $\pi_{k}$. In our adaptive algorithm for instance, $C^{\star}_{k}$ will intuitively grow as $\pi_{k}$ gets closer to $\pi_{\star}$. Note that the condition $C^{\star}_{k}$ is less restrictive than the uniform coverage assumption used in adaptive ; scrm . A similar inequality on the variance of importance weights was used as an assumption to obtain fast rates in scrm . Here, we prove such result and characterize when it is possible to obtain. For instance, this inequality cannot hold when the costs are constant. Finally, we used $\Delta_{u}$ and $C^{\star}_{k}$ for the sake of interpretability, and give a sharper, yet less interpretable condition in Sec. E.3.
Lemma 4.3 shows that the expansion of $L$ is controlled by a difference of risks. Smaller $\gamma_{k}$ sharpens this inequality, which is the case for problems with distinguishable optimal actions and near optimal logging policies $\pi_{k}$. Plugging this inequality in Proposition 4.2 then yields a recursive control on the sequence $\left(R(\pi_{k})\right)_{k\geq 0}$ (see Cor. E.5). Such a recursive relation can then be transformed in to an explicit accelerated convergence rate for the output of Algorithm 2 at the cost of the following mild assumption, where we suppose that our learning algorithm outputs for all $k\geq 0$, $\pi_{k}$ such that:
$\forall k\geq 0\,,\quad\gamma_{k}\leq\gamma<+\infty.$
This holds when $C^{\star}_{k}\geq C^{\star}>0$, which is a reasonable assumption as we expect $\pi_{0}$ to cover optimal actions and $\pi_{k},k\geq 1$ to be closer to $\pi_{\star}$. With this, we state our principal result below.
Theorem 4.4.
Let $n_{j}=m>0$ for all $j$. Then, for any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm 2 at time $k$ with parameter $\lambda_{m}:=\frac{1-\alpha}{8\gamma\sqrt{m}}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
$R(\pi_{k+1})-R(\pi^{\star})\leq\frac{C_{\alpha}}{\sqrt{m}(k+1)^{\alpha}},$
where $C_{\alpha}:=64\gamma\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{1-\alpha}$.
Thm. 4.4 is proven in Appendix E and provides an accelerated rate of $\mathcal{O}\left(\frac{1}{\sqrt{m}k^{\alpha}}\right)$ compared to the $\mathcal{O}\left(\frac{1}{\sqrt{mk}}\right)$ of Algorithm 1 as long as $\alpha>1/2$. Again, we recover the same rate than sakhi2024logarithmic in the batch setting (i.e. policy after one batch update: $k=1,m=n$) and improve the rate in the online setting (i.e. updating the behavior policy after each interaction: $k=n,m=1$) to $\mathcal{O}(1/n^{\alpha})$. Thm. 4.4 also works for any intermediate regime between batch and online settings, and suggests that considering several updates is beneficial to learn efficiently. Indeed, as long as $\alpha>1/2$, the number of updates $k$ will have a stronger influence than the size of the batch $m$. Note however that $C_{\alpha}$, the constant behind is deteriorating as $\alpha$ is close to $1$. Bigger values of $\lambda_{m}$ result in smaller constants but slower convergence rates, while smaller $\lambda_{m}$ trade a heavy constant for a fast convergence rate. There is then a practical balance to find between a fast convergence rate and reasonable values of $C_{\alpha}$.
Finally, to tackle a large-range of real-life scenarios, where $n_{k}$ may vary through time, we provide additional convergence rates Algorithm 2 encompassing this setting in Cor. 4.5.
Corollary 4.5.
Let $\beta_{1}:=\sup_{k\geq 0}\frac{k+1}{N_{k}}\leq 1$. Assume there exists $\beta_{2}\in\mathbb{R}^{+}$ such that for all $k$, $n_{k}\leq\beta_{2}$. For any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm 2 at time $k$ with parameter $\lambda=\frac{1}{1+2^{2+\alpha}\gamma\beta_{1}\beta_{2}B(\alpha)}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
$R(\pi_{k})-R(\pi^{\star})\leq\frac{C_{\alpha}}{(k+1)^{\alpha}},$
where $C_{\alpha}\leq\max\left(1,\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda^{2}\gamma\beta_{2}B(\alpha)}\right)$, and $B(\alpha)=\frac{1}{1-\alpha}$.
Observe that the convergence rate given in Cor. 4.5 is weaker than that of Thm. 4.4, as it does not directly depend on the batch size $n_{j}$. This relaxation is necessary to accommodate adversarial scenarios where $n_{j}$ can be as small as 1. For instance, consider $n\geq 2$ and a sequence such that for all $j\geq 1$, $n_{2j-1}=m$ and $n_{2j}=1$. In this case, for all $k\geq 0$, we have $N_{2k-1}=k(m+1)-1$ and $N_{2k}=k(m+1)$. Thus, for all $k$, it holds that $\frac{k}{N_{k}}\leq\frac{2}{m}$, and the assumptions of Cor. 4.5 are satisfied. Importantly, Cor. 4.5 matches the optimal convergence rate of Thm. 4.4 in the fully online regime (i.e., $k=n$, $n_{j}=1$). Having established that Algorithm 2 enjoys provably accelerated convergence rates, we turn to empirical validation to assess whether this theoretical advantage is reflected in practice.
5 Experiments
General Setup. Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work (dudik14doubly, ; swaminathan2015batch, ). Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies sakhi2023pac ; aouali23a , these policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost. Appendix F details the experimental design and obtained results.
Effect of Multiple Deployments. Unlike prior batch approaches, our adaptive algorithms support multiple updates of the behavior policy $\pi_{0}$, while seamlessly reducing to the batch case when $k=1$. Using the MNIST and CIFAR100 datasets, we evaluate how the number of intermediate policy updates $k$ affects the performance of the final learned policy $\pi$. Starting from an approximately uniform policy $\pi_{0}$, we run both adaLS (Alg 1) and the improved adaAdjLS (Alg 2). For each $k\in\{1,5,10,100\}$, we fix the total number of interactions to $N_{k}=N$, resulting in $m=N/k$ interactions per deployment. The case $k=1$ corresponds to the static setting studied in sakhi2024logarithmic , where all data is collected using $\pi_{0}$. For $k>1$, each intermediate policy $\pi_{j}$ interacts with $m$ examples, collectively producing $N$ adaptively collected samples. We set $\lambda\sim\mathcal{O}(1/\sqrt{m})$, following our theoretical guidance. Table 1 summarizes the results: increasing $k$ consistently improves final policy performance, with diminishing returns as $\pi$ nears optimality (e.g., MNIST at $k=10$ and $100$). The gains from adaptivity are especially notable on CIFAR100, where the larger action space ($|\mathcal{A}|=100$) makes uniform policies particularly suboptimal, thus highlighting the importance of the adaptive updates. Finally, our results demonstrate the empirical advantage of adaAdjLS, validating the need for our refined estimator. Across all settings, and especially the more challenging ones, adaAdjLS consistently outperforms adaLS, providing strong empirical evidence that the adjustment is essential for achieving accelerated convergence.
Table 1: $R(\pi)$ while fixing $N_{k}=N$, varying $k$ in MNIST and CIFAR100 (lower is better).
Dataset $k$ adaLS (Alg 1) adaAdjLS (Alg 2)
MNIST $1$ $-0.674$ $-0.678$
$5$ $-0.758$ $-0.766$
$10$ $-0.782$ $\mathbf{-0.791}$
$100$ $-0.781$ $\mathbf{-0.789}$
CIFAR100 $1$ $-0.197$ $-0.207$
$5$ $-0.344$ $-0.380$
$10$ $-0.406$ $-0.455$
$100$ $-0.508$ $\mathbf{-0.558}$
Figure 1: $R(\pi_{k})$ at each intermediate step, for $k=10$.
Benchmarks. Having demonstrated the benefits of intermediate updates and the effectiveness of our proposed adaAdjLS algorithm, we now compare our methods against the recently introduced SCRM approach scrm , which extends the CRM principle of swaminathan2015batch to the adaptive setting. We evaluate all methods on MNIST, FashionMNIST, EMNIST and CIFAR100 of different action space sizes, fixing the total number of interactions to $N$, setting $k=10$, and using theoretically recommended values for $\lambda$ for a fair comparison. adaLS and adaAdjLS divide the interaction budget uniformly across rounds ($n_{j}=N/k$), reusing all previously collected data to update policies. In contrast, SCRM employs exponentially increasing interaction sizes ($n_{j}=n_{0}\cdot 2^{j}$ with $n_{0}=\lceil N/2^{k}\rceil$) and updates policies using only the most recent batch, rather than the cumulative data ($n_{j}$ instead of $N_{j}$). Each method is run across six random seeds, and the results are summarized in Fig. 1. We find that SCRM tends to be overly conservative, particularly in early rounds where the number of interactions is small. While it does improve upon the initial behavior policy $\pi_{0}$ over time, it lags significantly behind both of our adaptive methods. Although it narrows the gap with adaLS on CIFAR100, it remains clearly outperformed—especially by adaAdjLS, which consistently achieves superior results across all datasets. These findings further confirm the empirical advantages of our approach. Additional experimental details and results can be found in Appendix F.
6 Conclusion
In this work, we extended the Logarithmic Smoothing learning principle to the adaptive setting. We derived two novel adaptive algorithms for off-policy learning and established convergence guarantees to their respective optima. One of the algorithms enjoys accelerated convergence under mild assumptions and demonstrated improved empirical performance. These theoretical and practical benefits highlight the potential of combining online PAC-Bayes techniques with adaptive contextual bandits. Importantly, online PAC-Bayes methods do not rely on the i.i.d. assumption, opening promising directions for future research. In particular, relaxing the assumptions of a fixed context distribution and i.i.d. contexts would allow us to address distribution shift—where the context or cost distribution evolves during training. This setting better reflects real-world applications, where user behavior and responses naturally change over time.
Acknowledgments and Disclosure of Funding
M. Haddouche is supported by the European Union (ERC grant DYNASTY 101039676). The French government partly funded this work under the management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI).
Appendix A Limitations
This work develops theoretically grounded and practical learning approaches for the adaptive contextual bandit setting, where the decision-maker can dynamically adjust behavior policies to collect higher-quality data. Our theoretical analysis generalizes the static (batch) setting and highlights the advantages of multiple updates and deployments. While this represents a clear improvement over the batch setting, the proposed algorithm is still reliant on the initial behavior policy and may not converge to the optimal policy if the starting policy fails to explore optimal actions. Another limitation arises from the offline contextual bandit framework itself, which assumes i.i.d. observations. Although this assumption may be valid in simple scenarios, it becomes inadequate in real-world applications where user interests evolve over time. Extending our results to settings with distribution shifts, for example, could be achieved using similar proof techniques and would present an important avenue for future research.
Appendix B Broader impact statement
We contribute to the development of principled and practical pessimistic algorithms for the adaptive off-policy contextual bandit setting. These algorithms improve the robustness of decision-making by reducing unnecessary exploration, accelerating convergence to optimal policies, and prioritizing safety while mitigating uncertainty-related risks. By incorporating pessimistic strategies, our approach introduces a conservative bias that can lead to better outcomes in high-stakes environments where the cost of mistakes is significant. While our framework and algorithms have broad potential applicability, their social impact will ultimately depend on the specific domains in which they are deployed.
Appendix C Extended Related Work
Off-Policy Learning.
Off-Policy Learning in the contextual bandit setting, or counterfactual risk minimization [47] is a widely adopted framework for learning from past interactions. In most real world applications, large datasets summarizing past interactions are often available, allowing agents to improve their policies offline. In this setting, the goal is to find the optimal policy using previously collected interactions. Two primary approaches exist for learning from this data. The direct method [43] focuses on modeling the reward and subsequently derives a straightforward policy. This approach is well-understood, and the quality of the learned policy depends entirely on our ability to accurately model the reward. The other approach directly learns a policy from the data by minimizing an importance weighting estimator [15, 6] of the expected cost. The research community has predominantly focused on this second approach, as it presents significant theoretical challenges and offers greater generality by not relying on any reward modeling. At the core of directly learning a policy is the development of tailored off-policy estimators with improved concentration properties [6, 3, 35, 42], combined with tight generalization bounds allowing efficient learning [47, 41, 42]. The used methods are motivated from the pessimism principle [26] that was proven to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens [31] proved efficient, improving the seminal work of [47]. This direction was further explored by [41, 3, 18, 42], reducing assumptions and improving the pessimistic principle, with latest developments [42] resulting in state of the art pessimistic approaches. We further develop this research direction and generalize it to the adaptive setting, where an online algorithm adapts the behavior policy to the environment to collect higher quality interactions.
Adaptive Off-Policy Learning.
Adaptive policy learning builds upon off-policy learning by enabling the decision-maker to refine the initial behavior policy and gather higher quality interactions with each new deployment. This approach is prevalent in practice as decision systems are continuously updated; however, its theoretical understanding remains limited and underexplored. Despite its practical significance, few works have addressed this question. [49] and subsequently [5] investigated learning from adaptively collected data in a purely online setting, where policies are updated after each step. In contrast to these online methods, our work considers updates occurring after the arrival of batches of arbitrary sizes. This better models real-world scenarios and provides a framework that interpolates between the fully static and fully online regimes. Our proposed algorithms achieve the convergence rate of the batch setting and accelerate it once we can refine our behavior policy. Furthermore, the analysis in [49, 5] relies on strong assumptions. Specifically, [49] employs a doubly robust estimator and provides regret guarantees but assumes uniform coverage to control variance. Under the same uniform coverage assumption, [5] introduces a novel maximal inequality and subsequently derives fast rate regret guarantees under a margin condition similar to ours. However, their results are limited to the online setting and also require uniform coverage, which can decay slowly over time. Recently, [48] extended Counterfactual Risk Minimization to the adaptive setting, allowing for intermediate policy deployments to facilitate the learning of better policies. Their algorithm achieves accelerated convergence rates under a Hölderian assumption but necessitates bounded importance weights, which, in the discrete action space case, is equivalent to the uniform coverage assumption. Additionally, their proposed algorithm only functions with exponentially increasing batch sizes, which is unrealistic, and only utilizes the last batch of interactions instead of the entire dataset. These limitations primarily stem from the theoretical tools employed, rendering the algorithm impractical. We analyze this problem from a PAC-Bayesian perspective and present the first algorithms that interpolate between full batch and full online regimes, matching the rates of both and operating under the coverage of the optimal actions alone.
PAC-Bayes learning.
PAC-Bayes learning is a recent branch of learning theory which emerged in the late 90s via the seminal work of [45, 32, 33, 34] and later pursued by [7, 8]. Modern surveys are available to describe the various advances in the field [19, 1]. PAC-Bayes provide generalization bounds involving a complexity term (often a KL divergence), inspired here from the Bayesian learning paradigm of designing a posterior knowledge of the learning problem based on both training data and a prior knowledge of the considered situation. On the contrary, PAC-Bayes, while inspired from the Bayesian philosophy, relies historically on tools from information theory. This general approach benefits from additional flexibility as PAC-Bayes can be linked and applied to Bayesian learning (see 19) but also blurs the notion of prior and posterior distributions, now independent of the fundamental Bayes formula. This flexibility allowed PAC-Bayes to reach various sub-fields of learning theory: optimization dynamics of learning algorithms [36, 22, 2], reinforcement learning [17], online learning [30, 20], contrastive learning [37], generative models [9] or federated learning [40, 27].
Appendix D Proofs of Section 3
In this whole section, we assume that all datasets $S_{k}$ have the same cardinal $n_{k}=m$. This section contains the proof of Sec. 3.2 and Algorithm 1.
D.1 Proof of Theorem 3.2
We start by restating the theorem of interest:
[…]
Proof of Sec. 3.2.
For any $k\geq 0$, let:
$f_{k}(S,\theta):=\sum_{j=0}^{k}\underbrace{\sum_{i\in I_{k}}-\log\left(1-\lambda R(d_{\theta})\right)+\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)}_{:=Y_{j}(S,\theta)}.$
Now for a given posterior $Q$ and a prior $P$, we properly apply the change of measure inequality for any $m$:
$\mathbb{E}_{\theta\sim Q}\left[f_{k}(S,\theta)\right]\leq\operatorname{KL}(Q,P)+\log\left(\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\right).$ (4)
We need to control the exponential random variable. To do so we prove the following lemma:
Lemma D.1.
The sequence $(M_{k}:=\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))_{k\geq 0}$ is a non-negative martingale.
Proof.
We fix $k\geq 0$. We show that $\mathbb{E}_{k-1}[M_{k}]=M_{k-1}$.
$\mathbb{E}_{k-1}[M_{k}]$ $=\mathbb{E}_{k-1}\left[\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta)\right]$
$=\mathbb{E}_{\theta\sim P}\left[\mathbb{E}_{k-1}\left[\exp\left(f_{k-1}(S,\theta)\right)\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}}{1-\lambda R(d_{\theta})}\right)\right)\right]\right]$
$=\mathbb{E}_{\theta\sim P}\left[\exp\left(f_{k-1}(S,\theta)\right)\mathbb{E}_{k-1}\left[\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}}{1-\lambda R(d_{\theta})}\right)\right)\right]\right].$
The second line comes from the application of Fubini as $P$ is data-independent and last line already demonstrated. Now we have for any $\theta\in\Theta$:
$\mathbb{E}_{k-1}\left[\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}}{1-\lambda R(d_{\theta})}\right)\right)\right]$ $=\mathbb{E}_{k-1}\left[\Pi_{i\in I_{k}}\frac{1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}}{1-\lambda R(d_{\theta})}\right]$
$=\Pi_{i\in I_{k}}\frac{1-\lambda\mathbb{E}_{k-1}\left[\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}\right]}{1-\lambda R(d_{\theta})},$
these computations holding as we exploited the assumption that, conditionally to $\mathcal{F}_{k-1}$, $S_{k}=\left((x_{i},a_{i},c_{i})\right)_{i\in I_{k}}$ is constituted of mutually independent random variables (thus an expectation of product equals a product of expectations). Now, recall that $\pi_{k}$ is $\mathcal{F}_{k-1}$-measurable and that for all $i\in I_{k},a_{i}\sim\pi_{k}(\cdot\mid x_{i})$. We then remark that, for any $\theta$, $i\in I_{k}$:
$\mathbb{E}_{k-1}\left[\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}\right]$ $=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left[\frac{d_{\theta}(a^{\prime}_{i}\mid x^{\prime}_{i})c^{\prime}_{i}}{\pi_{k}(a^{\prime}_{i}\mid x^{\prime}_{i})}\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\left[\frac{d_{\theta}(a^{\prime}_{i}\mid x^{\prime}_{i})}{\pi_{k}(a^{\prime}_{i}\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left(c^{\prime}_{i}\right)\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\left[c\left(x^{\prime}_{i},d_{\theta}(x^{\prime}_{i})\right)\right]$
$=R(d_{\theta})$
Thus $\mathbb{E}_{k-1}[M_{k}]=M_{k-1}$, this concludes the lemma’s proof. ∎
Now; given a martingale is also a supermartingale, we can apply Ville’s inequality which implies that with probability at least $1-\delta$, for any $k\geq 0$:
$\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\leq\frac{1}{\delta}.$
Plugging this back into Eq. 4 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any posterior $Q$:
$\sum_{j=0}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda R(d_{\theta})\right)+\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}^{0}(a_{i}\mid x_{i})}\right)\right]\leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
Now remark that for any $j\geq 0,i\in I_{j},\theta$ because $d_{\theta}(a,x)=\mathds{1}\{f_{\theta}(x)=a\}\in\{0,1\}$, we have for any $(a,x)$:
$\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)=d_{\theta}(a,x)\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right).$
We then have:
$\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{j}}-\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)\right]$ $=-\sum_{j=0}^{k}\sum_{i\in I_{j}}\pi_{Q}(a_{i}\mid x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)$
$=(k+1)m\lambda\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q}).$
We also use the convexity of $x\rightarrow-\log(1-\lambda x)$ to obtain:
$\sum_{j=0}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{j}}-\log\left(1-\lambda R(d_{\theta})\right)\right]$ $=m(k+1)\mathbb{E}_{\theta\sim Q}\left[-\log\left(1-\lambda R(d_{\theta})\right)\right]$
$\geq-m(k+1)\log\left(1-\lambda R(\pi_{Q})\right).$
Finally, mixing all those calculations together, dividing by $\lambda(k+1)m$ and applying the increasing function $\psi_{\lambda}(x):=\frac{1}{\lambda}(1-\exp(-\lambda x)$ gives:
$R(\pi_{Q})\leq\psi_{\lambda}\left(\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda(k+1)m}\right).$
Finally, using that $\psi_{\lambda}(x)\leq x$ concludes the proof. ∎
D.2 Proof of Proposition 1
We prove the convergence guarantee of Algorithm 1.
[…]
Proof.
We fix $k\geq 0$, an adaptive algorithm Alg and $P$ a data-free prior. To highlight the dependency on $\lambda$, we denote in this proof $\mathcal{S}_{j}=\mathcal{S}_{\lambda,j}$ Our goal is to apply the change of measure inequality on $\mathcal{H}$ to a specific function $f_{k}$. We define this function below, for any sample $S$ and any predictor $\theta$
$f_{k}(S,\theta):=\sum_{j=0}^{k}\underbrace{\sum_{i\in I_{k}}-\lambda R(d_{\theta})-\lambda^{2}\mathcal{S}_{\lambda,j}\left(d_{\theta}\right)-\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)}_{:=Y_{j}(S,\theta)},$
where $\mathcal{S}_{\lambda,j}(\pi):=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\pi(a\mid x)c^{2}}{\pi_{j}^{2}(a\mid x)-\lambda\pi_{j}(a\mid x)c}\right]$ Now for a given posterior $Q$ we apply the change of measure inequality for any $m$:
$\sum_{j=0}^{k}\mathbb{E}_{\theta\sim Q}\left[Y_{j}(S,\theta)\right]=\mathbb{E}_{\theta\sim Q}\left[f_{k}(S,\theta)\right]\leq\operatorname{KL}(Q,P)+\log\left(\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\right).$ (5)
We need to control the exponential random variable. To do so we prove the following lemma:
Lemma D.2.
The sequence $(M_{k}:=\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))_{k\geq 0}$ is a non-negative supermartingale.
Proof.
We fix $k\geq 0$ and we recall that for any $k$, $P$ is data free. We show that $\mathbb{E}_{k-1}[M_{k}]\leq M_{k-1}$.
$\mathbb{E}_{k-1}[M_{k}]$ $=\mathbb{E}_{k-1}\left[\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta)\right]$
$=\mathbb{E}_{\theta\sim P}\left[\mathbb{E}_{k-1}\exp(f_{k}(S,\theta)\right]$
$=\mathbb{E}_{\theta\sim P}\left[\mathbb{E}_{k-1}\left[\exp\left(f_{k-1}(S,\theta)\right)\exp\left(Y_{j}(S,\theta\right)\right]\right]$
$=\mathbb{E}_{\theta\sim P}\left[\exp\left(f_{k-1}(S,\theta)\right)\mathbb{E}_{k-1}\left[\exp\left(Y_{j}(S,\theta\right)\right]\right].$
The second line holding, as $P$ is data free, by Fubini.
Now we note $c=c(a,x)$ for the sake of conciseness. For any $K,\theta\in\mathbb{R}^{d}$, using the fact that conditionally to $\mathcal{F}_{k-1}$, all the data in $\mathcal{S}_{k}$ are independent we have:
$\mathbb{E}_{k-1}\left[\exp\left(Y_{j}(S,\theta\right)\right]$ $=\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-\lambda\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)\right)\right)\mathbb{E}_{k-1}\left[\frac{1}{1-\lambda\frac{d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)}}\right]$
Using $\log(x)\leq x-1$ yields
$\leq\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-\lambda\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)\right)+\mathbb{E}_{k-1}\left[\frac{1}{1-\lambda\frac{d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)}}\right]-1\right)$
$=\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-\lambda\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)\right)+\mathbb{E}_{k-1}\left[\frac{\lambda d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)-\lambda d_{\theta}(a\mid x)c}\right]\right)$
Using that $d_{\theta}$ is binary gives the equality $\frac{\lambda d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)-\lambda d_{\theta}(a\mid x)c}=\frac{\lambda d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)-\lambda c}$, thus
$=\prod_{i\in I_{k}}\left(\exp\left(-\lambda\left(R\left(d_{\theta}\right)-\lambda\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)\right)+\mathbb{E}_{k-1}\left[\frac{\lambda d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)-\lambda c}\right]\right)\right)$
$=\prod_{i\in I_{K}}\left(\exp\left(-\lambda^{2}\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)+\mathbb{E}_{k-1}\left[\frac{\lambda d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)-\lambda c}-\lambda\frac{d_{\theta}(a\mid x)c}{\pi_{k}(a\mid x)}\right]\right)\right)$
$\leq\prod_{i\in I_{k}}\left(\exp\left(-\lambda^{2}\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)+\lambda^{2}\mathcal{S}_{\lambda,k}\left(d_{\theta}\right)\right)\right)$
$\leq 1.$
This allows concluding that $\mathbb{E}_{k-1}[M_{k}]\leq M_{k-1}$, thus the proof. ∎
Now, we can apply Ville’s inequality which implies that with probability at least $1-\delta$, for any $K\geq 1$:
$\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\leq\frac{1}{\delta}.$
Plugging this back into Eq. 5 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any $Q$:
$\mathbb{E}_{\theta\sim Q}\left[\sum_{j=0}^{k}\sum_{i\in I_{k}}-\lambda R(d_{\theta})-\lambda\mathcal{S}_{\lambda,j}\left(d_{\theta}\right)-\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)\right]\\ \leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
Then recall that
$\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q}):=-\frac{1}{(k+1)m\lambda}\sum_{j=1}^{k}\sum_{i\in I_{j}}\pi_{Q}(a_{i}\mid x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right).$
We have, using again that $d_{\theta}$ is binary for any $\theta$:
$\mathbb{E}_{\theta\sim Q}\left[\sum_{j=0}^{k}\sum_{i\in I_{k}}-\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{i}^{0}(a_{i}\mid x_{i})}\right)\right]$ $=-\sum_{j=0}^{k}\sum_{i\in I_{k}}\pi_{Q}(a_{i}\mid x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)$
$=(k+1)m\lambda\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q}).$
Mixing all those calculations together and dividing by $\lambda(k+1)m$ gives: with probability at least $1-\delta$, for any $k$, $Q$:
$\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q})\leq R(\pi_{Q})+\frac{\lambda}{k+1}\sum_{j=0}^{k}\mathcal{S}_{\lambda,j}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda(k+1)m}$ (6)
Now, we consider the optimal sequence, defined such that, $Q^{\star}:=\operatorname{argmin}_{Q}R(\mathrm{Q})$.
We also consider $\pi_{k+1}$ the output of Algorithm 1 time $k$ satisfying:
$\pi_{k+1}=\pi_{\hat{Q}_{k}},\quad\text{with}\;\hat{Q}_{k}:=\operatorname{argmin}_{Q}\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q})+\frac{\operatorname{KL}(Q,P)}{\lambda(k+1)m}$
Then using Sec. 3.2, and the definition of $\hat{Q}$ yields for any $\lambda$, data-free $P$ with probability at least $21-\delta/$ 2, for any $k$:
$R\left(\pi_{k+1}\right)$ $\leq\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{k+1})+\frac{\operatorname{KL}(\hat{Q}_{k},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
$\leq\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\star})+\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
Then, exploiting Eq. 6 with probability $21-\delta/$ 2 and taking a union bound yields, with probability at least $1-\delta$, for any $k$:
$R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{k+1}\sum_{j=0}^{k}\mathcal{S}_{\lambda,j}\left(\pi_{\star}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}.$ (7)
Finally, using that $\mathcal{S}_{\lambda,j}\leq\mathcal{S}_{j}$ for all $j$ concludes the proof. ∎
Appendix E Proofs of Section 4
E.1 Proof of Theorem 4.1
Proof.
For any $k\geq 0$, let:
$Y_{i,j}:=\log\left(1-\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{j}\right)\right)\right)-\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}-1\right)c_{i}\right)$
$f_{k}(S,\theta):=-\sum_{j=1}^{k}\sum_{i\in I_{j}}Y_{i,j}.$
Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality [12, 13] for any $k$:
$\mathbb{E}_{\theta\sim Q}\left[f_{k}(S,\theta)\right]\leq\operatorname{KL}(Q,P)+\log\left(\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\right).$ (8)
We need to control the exponential random variable. To do so we prove the following lemma:
Lemma E.1.
The sequence $(M_{k}:=\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))_{k\geq 0}$ is a non-negative martingale.
Proof.
We fix $k\geq 0$. We show that $\mathbb{E}_{k-1}[M_{k}]=M_{k-1}$.
$\mathbb{E}_{k-1}[M_{k}]$ $=\mathbb{E}_{k-1}\left[\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta)\right]$
$=\mathbb{E}_{\theta\sim P}\left[\mathbb{E}_{k-1}\left[\exp\left(f_{k-1}(S,\theta)\right)\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}}{1-\lambda(R(d_{\theta})-R(\pi_{k}))}\right)\right)\right]\right]$
$=\mathbb{E}_{\theta\sim P}\left[\exp\left(f_{k-1}(S,\theta)\right)\mathbb{E}_{k-1}\left[\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}}{1-\lambda(R(d_{\theta})-R(\pi_{k}))}\right)\right)\right]\right].$
The second line comes from the application of Fubini as $P$ is data-independent. Now we have for any $\theta\in\Theta$:
$\mathbb{E}_{k-1}\left[\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}}{1-\lambda(R(d_{\theta})-R(\pi_{k}))}\right)\right)\right]\\ =\mathbb{E}_{k-1}\left[\Pi_{i\in I_{k}}\frac{1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}}{1-\lambda(R(d_{\theta})-R(\pi_{k}))}\right]\\ =\Pi_{i\in I_{k}}\frac{1-\lambda\mathbb{E}_{k-1}\left[\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}\right]}{1-\lambda(R(d_{\theta})-R(\pi_{k}))},$
these computations holding as we exploited the assumption that, conditionally to $\mathcal{F}_{k-1}$, $S_{k}=\left((x_{i},a_{i},c_{i})\right)_{i\in I_{k}}$ is constituted of mutually independent random variables (thus an expectation of product equals a product of expectations). Now, recall that $\pi_{k}$ is $\mathcal{F}_{k-1}$-measurable and that for all $i\in I_{k},a_{i}\sim\pi_{k}(\cdot\mid x_{i})$. We then remark that, for any $\theta$, $i\in I_{k}$:
$\mathbb{E}_{k-1}\left[\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{k}(a_{i}\mid x_{i})}\right]$ $=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left[\frac{d_{\theta}(a^{\prime}_{i}\mid x^{\prime}_{i})c^{\prime}_{i}}{\pi_{k}(a^{\prime}_{i}\mid x^{\prime}_{i})}\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\left[\frac{d_{\theta}(a^{\prime}_{i}\mid x^{\prime}_{i})}{\pi_{k}(a^{\prime}_{i}\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left(c^{\prime}_{i}\right)\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{k}(\cdot\mid x^{\prime}_{i})}\left[c\left(x^{\prime}_{i},d_{\theta}(x^{\prime}_{i})\right)\right]$
$=R(d_{\theta})$
Similarly, notice that $\mathbb{E}_{k-1}[c_{i}]=R(\pi_{k})$.
Thus $\mathbb{E}_{k-1}[M_{k}]=M_{k-1}$, this concludes the lemma’s proof. ∎
Now; given a martingale is also a supermartingale, we can apply Ville’s inequality [14], which implies that with probability at least $1-\delta$, for any $k\geq 0$:
$\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\leq\frac{1}{\delta}.$
Plugging this into Eq. 8 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
$\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{j}}-\log\left(1-\lambda\left(R(d_{\theta})-R(\pi_{j})\right)\right)+\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{j}(a_{i}\mid x_{i})}-1\right)c_{i}\right)\right]\leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
Now remark that for any $j\geq 1,i\in I_{j},\theta$ because $d_{\theta}(a,x)=\mathds{1}\{f_{\theta}(x)=a\}\in\{0,1\}$, we have for any $(a,x)$:
$\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{j}(a_{i}\mid x_{i})}-1\right)c_{i}\right)=d_{\theta}(a_{i}|x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})(1+\lambda c_{i})}\right)-\log(1+\lambda c_{i}).$
We then have:
$\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{j}}-\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{j}(a_{i}\mid x_{i})}-1\right)c_{i}\right)\right]$ $=N_{k}\lambda\left(\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)\right).$
We also use the convexity of $x\rightarrow-\log(1-\lambda x)$ to obtain:
$\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{j}}-\log\left(1-\lambda\left(R(d_{\theta})-R(\pi_{j})\right)\right)\right]$ $=-\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[n_{j}\log\left(1-\lambda\left(R(d_{\theta})-R(\pi_{j})\right)\right)\right]$
$=-N_{k}\sum_{j=1}^{k}\mathbb{E}_{\theta\sim Q}\left[\frac{n_{j}}{N_{k}}\log\left(1-\lambda\left(R(d_{\theta})-R(\pi_{j})\right)\right)\right]$
$\geq-N_{k}\log\left(1-\lambda\left(R(\pi_{Q})-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R(\pi_{j})\right)\right).$
Finally, mixing all those calculations together, dividing by $\lambda N_{k}$ and applying the increasing function $\psi_{\lambda}(x):=\frac{1}{\lambda}(1-\exp(-\lambda x)$ gives:
$R(\pi_{Q})\leq\psi_{\lambda}\left(\hat{R}_{0,k}^{\lambda}(\pi_{Q})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}}\right).$
Finally, using that $\psi_{\lambda}(x)\leq x$ concludes the proof. ∎
E.2 Proof of Proposition 4.2
We need a preliminary proposition to prove Proposition 4.2.
After we proved our upper bound, we need another theorem to analyze the algorithm.
Proposition E.2.
For any adaptive algorithm Alg, data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)$ $\leq R\left(\pi_{Q}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)+\frac{\lambda}{1-\lambda}\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{Q},\pi_{j}\right)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}}$
with $L\left(\pi_{Q},\pi_{j}\right)=\mathbb{E}_{x\sim\nu}\left[\mathbb{E}_{\pi_{j}}\left[c^{2}\right]+\mathbb{E}_{\pi_{Q}}\left[c^{2}\left(\frac{1}{\pi_{j}(a|x)}-2\right)\right]\right].$
Proof.
For any $k\geq 0$, and $\lambda\in(0,1)$, let:
$Y_{i,j}:=\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}-1\right)c_{i}\right)+\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{j}\right)+\lambda M_{\lambda}(d_{\theta},\pi_{j})\right)$
$f_{k}(S,\theta):=-\sum_{j=1}^{k}\sum_{i\in I_{j}}Y_{i,j}.$
with $M_{\lambda}(d_{\theta},\pi_{j})=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}}{1-\lambda\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)c}\right].$ Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality [12, 13] for any $k$:
$\mathbb{E}_{\theta\sim Q}\left[f_{k}(S,\theta)\right]\leq\operatorname{KL}(Q,P)+\log\left(\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\right).$ (9)
We also need to control the exponential random variable. We prove the following lemma:
Lemma E.3.
The sequence $(M_{k}:=\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta)))_{k\geq 0}$ is a non-negative supermartingale.
Proof.
We fix $k\geq 0$. We show that $\mathbb{E}_{k-1}[M_{k}]\leq M_{k-1}$.
Let $L_{k}=\mathbb{E}_{k-1}\left[\exp\left(\sum_{i\in I_{k}}\log\left(\frac{1}{1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{k}(a_{i}\mid x_{i})}-1\right)c_{i}}\right)-\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{k}\right)+\lambda M_{\lambda}(d_{\theta},\pi_{k})\right)\right)\right].$ We have:
$\mathbb{E}_{k-1}[M_{k}]$ $=\mathbb{E}_{k-1}\left[\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta)\right]$
$=\mathbb{E}_{\theta\sim P}\left[\exp\left(f_{k-1}(S,\theta)\right)L_{k}\right],$
where the last line comes from Fubini’s theorem as $P$ is data-independent. Let us prove that $L_{k}$ is smaller than $1$. We have:
$L_{k}$ $=\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{k}\right)+\lambda M_{\lambda}\left(d_{\theta},\pi_{k}\right)\right)\right)\mathbb{E}_{k-1}\left[\frac{1}{1-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}\right]$
Using $\log(x)\leq x-1$ yields
$L_{k}$ $\leq\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{k}\right)+\lambda M_{\lambda}\left(d_{\theta},\pi_{k}\right)\right)+\mathbb{E}_{k-1}\left[\frac{1}{1-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}\right]-1\right)$
$=\prod_{i\in I_{k}}\exp\left(-\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{k}\right)+\lambda M_{\lambda}\left(d_{\theta},\pi_{k}\right)\right)+\mathbb{E}_{k-1}\left[\frac{\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}{1-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}\right]\right)$
$=\prod_{i\in I_{k}}\exp\left(-\lambda^{2}M_{\lambda}\left(d_{\theta},\pi_{k}\right)+\mathbb{E}_{k-1}\left[\frac{\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}{1-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c\right]\right)$
$=\prod_{i\in I_{k}}\exp\left(-\lambda^{2}M_{\lambda}\left(d_{\theta},\pi_{k}\right)+\mathbb{E}_{k-1}\left[\frac{\lambda^{2}(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)^{2}c^{2}}{1-\lambda(\frac{d_{\theta}(a\mid x)}{\pi_{k}(a\mid x)}-1)c}\right]\right)$
$=\prod_{i\in I_{k}}\exp\left(-\lambda^{2}M_{\lambda}\left(d_{\theta},\pi_{k}\right)+\lambda^{2}M_{\lambda}\left(d_{\theta},\pi_{k}\right)\right)$
$=1.$
Thus $\mathbb{E}_{k-1}[M_{k}]\leq M_{k-1}$, this concludes the lemma’s proof. ∎
We can apply Ville’s inequality [14] which implies that with probability at least $1-\delta$, for any $k\geq 0$:
$\mathbb{E}_{\theta\sim P}\exp(f_{k}(S,\theta))\leq\frac{1}{\delta}.$
Plugging this into Eq. 9 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
$-\sum_{j=0}^{k}\sum_{i\in I_{j}}\mathbb{E}_{\theta\sim Q}\left[Y_{i,j}\right]\leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
with $Y_{i,j}=\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}-1\right)c_{i}\right)+\lambda\left(R\left(d_{\theta}\right)-R\left(\pi_{j}\right)+\lambda M_{\lambda}(d_{\theta},\pi_{j})\right)$.
Now remark that for any $j\geq 1,i\in I_{j},\theta$ because $d_{\theta}(a,x)=\mathds{1}\{f_{\theta}(x)=a\}\in\{0,1\}$, we have for any $(a,x)$:
$\log\left(1-\lambda\left(\frac{d_{\theta}(a_{i}\mid x_{i})}{\pi_{j}(a_{i}\mid x_{i})}-1\right)c_{i}\right)=d_{\theta}(a_{i}|x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})(1+\lambda c_{i})}\right)-\log(1+\lambda c_{i}).$
We thus obtain by re-organising the terms and dividing by $\lambda N_{k}$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)$ $\leq R\left(\pi_{Q}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)+\lambda\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}\mathbb{E}_{\theta\sim Q}\left[M_{\lambda}\left(d_{\theta},\pi_{j}\right)\right]+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}}.$
We proceed then by bounding $\mathbb{E}_{\theta\sim Q}\left[M_{\lambda}\left(d_{\theta},\pi_{j}\right)\right]$:
$\mathbb{E}_{\theta\sim Q}\left[M_{\lambda}\left(d_{\theta},\pi_{j}\right)\right]$ $=\mathbb{E}_{\theta\sim Q}\left[M_{\lambda}\left(d_{\theta},\pi_{j}\right)\right]$
$=\mathbb{E}_{\theta\sim Q}\left[\mathbb{E}_{\pi_{j}}\left[\frac{\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}}{1-\lambda\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)c}\right]\right]$
$\leq\frac{1}{1-\lambda}\mathbb{E}_{\theta\sim Q}\left[\mathbb{E}_{\pi_{j}}\left[\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}\right]\right]$
$=\frac{1}{1-\lambda}\mathbb{E}_{\pi_{j}}\left[\mathbb{E}_{\theta\sim Q}\left[\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}\right]\right]$
$=\frac{1}{1-\lambda}\mathbb{E}_{\pi_{j}}\left[\mathbb{E}_{\theta\sim Q}\left[d_{\theta}(a|x)\left(\frac{1}{\pi_{j}(a|x)}-1\right)^{2}c^{2}+(1-d_{\theta}(a|x))c^{2}\right]\right]$
$=\frac{1}{1-\lambda}\mathbb{E}_{\pi_{j}}\left[\pi_{Q}(a|x)\left(\frac{1}{\pi_{j}(a|x)}-1\right)^{2}c^{2}+(1-\pi_{Q}(a|x))c^{2}\right]$
$=\frac{1}{1-\lambda}\mathbb{E}_{\pi_{j}}\left[\pi_{Q}(a|x)\left(\frac{1}{\pi_{j}(a|x)}-1\right)^{2}c^{2}+(1-\pi_{Q}(a|x))c^{2}\right]$
$=\mathbb{E}_{\pi_{j}}\left[c^{2}\right]+\mathbb{E}_{\pi_{Q}}\left[c^{2}\left(\frac{1}{\pi_{j}(a|x)}-2\right)\right]=L(\pi_{Q},\pi_{j}).$
We used several times above the property $d_{\theta}\in\{0,1\}$ to transform the analytical expression of the random variable with $\mathbb{E}_{\pi_{j}}$. This gives our result and ends the proof. ∎
Now we are able to prove our main result. Let us recall it below for conciseness.
Theorem E.4.
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have, denoting by $\pi_{j}$ the output of $\mathcal{A}$ at time $j$, for all $k\geq 0$:
$0\leq R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{1-\lambda}\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}}$
with $\pi_{k+1}$ defined as the output of the adaptive algorithm Algorithm 2 at time $k$.
Proof.
We know that by Thm. 4.1, for any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $21-\delta/$ 2 over $S$, we have for all $k\geq 0$:
$R\left(\pi_{k+1}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$ $\leq\psi_{\lambda}\left(\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{k+1})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q_{k+1},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}}\right)$
$\leq\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{k+1})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q_{k+1},P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}},$
as $\psi_{\lambda}(x)\leq x$. Then, by definition, $\hat{Q}_{k+1}=\operatorname{argmin}_{Q}\left\{\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda N_{k}}\right\}$, thus:
$\forall k\geq 0,\quad R\left(\pi_{k+1}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$ $\leq\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\star})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}}.$
Now, we exploit Proposition E.2, with probability at least $21-\delta/$ 2 over $S$, we have $\forall k\geq 0$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\star})+\hat{C}_{0,k}(\lambda)$ $\leq R\left(\pi_{\star}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)+\frac{\lambda}{1-\lambda}\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)+\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}}.$
Now, taking an union bound and combining the two previous results gives, with probability at least $1-\delta$ over $S$, $\forall k\geq 0$:
$R\left(\pi_{k+1}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$ $\leq R\left(\pi_{\star}\right)-\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)+\frac{\lambda}{1-\lambda}\sum_{j=1}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)$
$\quad+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}}.$
Re-organising the terms conclude the proof. ∎
E.3 Proof of Lemma 4.3
Proof.
We are looking for $\gamma_{k}>0$ such that:
$\frac{L(\pi_{\star},\pi_{k})}{R(\pi_{k})-R(\pi_{\star})}\leq\gamma_{k}.$
Let us start rewriting the suboptimality $R(\pi_{k})-R(\pi_{\star})$. We have:
$R(\pi_{k})-R(\pi_{\star})$ $=\mathbb{E}_{x}\left[\mathbb{E}_{\pi_{k}}\left[c(a,x)\right]-c(a_{\star}(x),x)\right]$
$=\mathbb{E}_{x}\left[\mathbb{E}_{\pi_{k}}\left[c(a,x)-c(a_{\star}(x),x)\right]\right]$
$=\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right].$
On the other hand, we have by definition of $L(\pi_{\star},\pi_{k})$:
$L(\pi_{\star},\pi_{k})$ $=\mathbb{E}_{x}\left[\mathbb{E}_{\pi_{k}}\left[\mathbb{E}_{p(\cdot|x,a)}\left[c^{2}\right]\right]+\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-2\right)\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]$
$=\mathbb{E}_{x}\left[\mathbb{E}_{\pi_{k}}\left[\mathbb{E}_{p(\cdot|x,a)}\left[c^{2}\right]-\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]+\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-1\right)\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]$
$=\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\left[\mathbb{E}_{p(\cdot|x,a)}\left[c^{2}\right]-\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]+\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-1\right)\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]$
$=\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\left[\sigma^{2}(a,x)-\sigma^{2}(a_{\star}(x),x)+c^{2}(a,x)-c^{2}(a_{\star}(x),x)\right]\right]$
$+\mathbb{E}_{x}\left[\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-1\right)\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right].$
As the costs are in $[0,1]$, we can bound the variance term by:
$\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\left[\sigma^{2}(a,x)-\sigma^{2}(a_{\star}(x),x)\right]\right]\leq\frac{1}{4}\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\right]=\frac{1}{4}\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right].$
We then use the definition of $\Delta_{x,a}$ to develop $c(a,x)^{2}$ and obtain:
$c(a,x)^{2}$ $=\left(\Delta_{a,x}+c(a_{\star}(x),x)\right)^{2}$
$=\Delta_{a,x}^{2}+2c(a_{\star}(x),x)\Delta_{x,a}+c(a_{\star}(x),x)^{2}\leq\Delta_{a,x}^{2}+c(a_{\star}(x),x)^{2}.$
Plugging these two results into our previous equation, we get:
$L(\pi_{\star},\pi_{k})$ $\leq\frac{1}{4}\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]+\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{x,a}^{2}\right]+\mathbb{E}_{x}\left[\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-1\right)\mathbb{E}_{p(\cdot|x,a_{\star}(x))}\left[c^{2}\right]\right]$
$\leq\frac{1}{4}\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]+\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{x,a}^{2}\right]+\mathbb{E}_{x}\left[\left(\frac{1}{\pi_{k}(a^{\star}(x)|x)}-1\right)\right]$
$\leq\frac{1}{4}\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]+\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{x,a}^{2}\right]+\frac{1}{C^{\star}_{k}}\mathbb{E}_{x}\left[1-\pi_{k}(a^{\star}(x)|x)\right].$
We then proceed at upper bound the quantity:
$\frac{L(\pi_{\star},\pi_{k})}{R(\pi_{k})-R(\pi_{\star})}$ $\leq\frac{\frac{1}{4}\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]+\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{x,a}^{2}\right]+\frac{1}{C^{\star}_{k}}\mathbb{E}_{x}\left[1-\pi_{k}(a^{\star}(x)|x)\right]}{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right]}$
$\leq 1+\left(\frac{1}{4}+\frac{1}{C_{k}^{\star}}\right)\frac{\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]}{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right]},$
because $\Delta_{x,a}^{2}\leq\Delta_{x,a}$ as $\Delta_{x,a}\leq 1$.
Now, we proceed to define:
$\bar{\Delta}_{k}=\frac{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right]}{\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]}.$
$\bar{\Delta}_{k}$ represents the expected sub-optimality incurred when policy $\pi_{k}$ selects a non-optimal action. A large $\bar{\Delta}_{k}$ means that the environment penalizes us more for not choosing an optimal action, while $\bar{\Delta}_{k}\approx 0$ means that the environment has multiple near-optimal actions. We can characterize $\gamma_{k}$ by this quantity, or loosen our bound and have a more interpretable condition. We choose the second approach. Let us suppose that there exists a $\Delta_{u}>0$ for a $u\in[0,1)$ such that:
$P_{x}\left(\min_{a\neq a^{\star}(x)}\Delta_{x,a}\geq\Delta_{u}\right)=1-u.$
Let us focus on the numerator of $\bar{\Delta}_{k}$. We have:
$P\left(\Delta_{x,a}\geq\Delta_{u}|a\neq a^{\star}(x)\right)\geq P_{x}\left(\min_{a\neq a^{\star}(x)}\Delta_{x,a}\geq\Delta_{u}\right)=1-u.$
Indeed, the event $E_{1}=\{\Delta_{x,a}\geq\Delta_{u}|a\neq a^{\star}(x)\}$ represents choosing a random action from the suboptimal set, and finding that $\Delta_{x,a}\geq\Delta_{u}$, while the event $E_{0}=\{\min_{a\neq a^{\star}(x)}\Delta_{x,a}\geq\Delta_{u}\}$ means that all non-optimal actions have a sub-optimality bigger than $\Delta_{u}$. We have $E_{0}\subset E_{1}$, and thus:
$P\left(\Delta_{x,a}\geq\Delta_{u}|a\neq a^{\star}(x)\right)\geq 1-u.$
Additionally, we have by the definition of conditional expectations:
$P\left(\Delta_{x,a}\geq\Delta_{u}|a\neq a^{\star}(x)\right)$ $=\frac{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\mathds{1}\left[\Delta_{a,x}>\Delta_{u}\right]\right]}{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\right]}$
$=\frac{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\mathds{1}\left[\Delta_{a,x}>\Delta_{u}\right]\right]}{\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]}\geq 1-u.$
Now, we write:
$\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right]$ $\geq\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\mathds{1}\left[\Delta_{a,x}>\Delta_{u}\right]\right]$
$\geq\Delta_{u}\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\mathds{1}\left[\Delta_{a,x}>\Delta_{u}\right]\right]$
$\geq\Delta_{u}(1-u)\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right].$
Finally, we have:
$\bar{\Delta}_{k}=\frac{\mathbb{E}_{x}\left[\sum_{a\neq a_{\star}(x)}\pi_{k}(a|x)\Delta_{a,x}\right]}{\mathbb{E}_{x}\left[1-\pi_{k}(a_{\star}(x)|x)\right]}\geq\Delta_{u}(1-u).$
and we finally obtain our result:
$\frac{L(\pi_{\star},\pi_{k})}{R(\pi_{k})-R(\pi_{\star})}$ $\leq 1+\left(\frac{1}{4}+\frac{1}{C_{k}^{\star}}\right)\frac{1}{\Delta_{u}(1-u)}=\gamma_{k}$
This ends the proof. ∎
E.4 Proof of Theorem 4.4
We first start with a key intermediary result, derived from Lemma 4.3 coupled with Proposition 4.2.
Corollary E.5.
Let $\pi_{k+1}$ be the output of Algorithm 2 at time $k$ with parameter $\lambda\in(0,1)$. For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$:
$0\leq D_{k+1}$ $\leq\frac{\lambda}{1-\lambda}\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}\gamma_{j}D_{j}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}},$
where $D_{k}=R(\pi_{k})-R(\pi^{\star})$.
We now proceed to prove Theorem 4.4.
Proof.
Assume that for all $j$, $n_{j}=m$, then $N_{k}=(k+1)m$ for all $k$. We re-use the notation $D_{k}=R(\pi_{k})-R(\pi^{\star})$ for all $k\geq 0$. We define
$\lambda_{m}$ $:=\frac{1-\alpha}{8\gamma\sqrt{m}}$
$C_{\alpha}$ $:=64\gamma\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{(1-\alpha)}.$
Now let us prove the result by strong recursion.
Initialisation
Take $k=0$. We have, by Cor. E.5:
$D_{1}$ $\leq\frac{\lambda}{1-\lambda}\gamma_{1}D_{0}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda m},$
$\leq\frac{\lambda}{1-\lambda}\gamma D_{0}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda m}$
Now we take $\lambda=\lambda_{m}$ and notice that $\lambda_{m}\leq 1/2$, thus $\frac{\lambda_{m}}{1-\lambda_{m}}\leq 2\lambda_{m}$:
$\leq\frac{1}{\sqrt{m}}\left(\frac{(1-\alpha)}{4}D_{0}+16\gamma\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{1-\alpha}\right),$
$\leq\frac{1}{\sqrt{m}}\left(\frac{(1-\alpha)}{4}D_{0}+16\gamma\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{1-\alpha}\right),$
Now as we have $\frac{1-\alpha}{4}D_{0}\leq\frac{1}{4}\leq\frac{C_{\alpha}}{2}$ we write:
$\leq\frac{1}{\sqrt{m}}\left(\frac{C_{\alpha}}{2}+\frac{C_{\alpha}}{4}\right),$
This then ensures that $D_{1}\leq\frac{C_{\alpha}}{\sqrt{m}}$ and prove the result for $k=1$
General case.
Assume the result true for all $1\leq j\leq k$. By Cor. E.5, we have for any $\lambda\leq 1/2$:
$D_{k+1}$ $\leq\frac{\lambda}{1-\lambda}\frac{1}{k+1}\sum_{j=0}^{k}\gamma_{j}D_{j}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m},$
$\leq 2\lambda\gamma\frac{1}{k+1}\sum_{j=0}^{k}D_{j}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda km},$
Now, we use the recursion assumption to get
$D_{k+1}$ $\leq 2\lambda\gamma\left(\frac{D_{0}}{k+1}+\frac{1}{(k+1)\sqrt{m}}\sum_{j=1}^{k}\frac{C_{\alpha}}{k^{\alpha}}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m},$
$\leq 2\lambda\gamma\left(\frac{D_{0}}{k+1}+\frac{C_{\alpha}B(\alpha)}{(k+1)\sqrt{m}}k^{1-\alpha}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
$\leq 2\lambda\gamma\left(\frac{D_{0}}{k+1}+\frac{C_{\alpha}B(\alpha)}{(k+1)^{\alpha}\sqrt{m}}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m},$
Now as $\lambda_{m}\leq 1/2$, take $\lambda=\lambda_{m}$, we have:
$D_{k+1}$ $\leq\frac{1}{\sqrt{m}}\left(\frac{D_{0}(1-\alpha)}{4(k+1)}+\frac{C_{\alpha}}{4(k+1)^{\alpha}\sqrt{m}}\right)+16\gamma\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{(1-\alpha)(k+1)\sqrt{m}},$
$\leq\frac{1}{\sqrt{m}(k+1)^{\alpha}}\left(\frac{D_{0}(1-\alpha)}{4}+\frac{C_{\alpha}}{4\sqrt{m}}+16\gamma\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{(1-\alpha)}\right),$
We use $C_{\alpha}\geq D_{0}$ and its expression to get:
$D_{k+1}$ $\leq\frac{1}{\sqrt{m}(k+1)^{\alpha}}\left(\frac{C_{\alpha}}{4}+\frac{C_{\alpha}}{4\sqrt{m}}+\frac{C_{\alpha}}{4}\right),$
We then have: $D_{k+1}\leq\frac{C_{\alpha}}{\sqrt{m}(k+1)^{\alpha}}$, this concludes the proof. ∎
E.5 Proof of Corollary 4.5
Proof of Cor. 4.5.
We re-use the notation $D_{k}=R(\pi_{k})-R(\pi^{\star})$ for all $k\geq 0$. We define, for any $\alpha\in[0,1)$,
$C_{\alpha}:=\max\left(D_{0},(1-\lambda)\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda^{2}\gamma^{*}\beta_{2}B(\alpha)}\right)\leq\max\left(1,\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda^{2}\gamma^{*}\beta_{2}B(\alpha)}\right),$
where $\lambda=\frac{1}{1+2^{2+\alpha}\gamma^{\star}\beta_{1}\beta_{2}B(\alpha)}$ Note that thanks to this definition, we have
$C_{\alpha}\frac{\lambda}{1-\lambda}\gamma^{*}\beta_{2}B(\alpha)\geq\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda}.$ (10)
Now let us prove the result by strong recursion.
Initialisation.
By definition of $C_{\alpha}$, we have $D_{0}\leq C_{\alpha}$.
General case.
For $k\geq 0$, Assume the result true for all $j\leq k$. By Cor. E.5, we have:
$D_{k+1}$ $\leq\frac{\lambda}{1-\lambda}\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}\gamma_{j}D_{j}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}},$
$\leq\frac{\lambda}{1-\lambda}\gamma^{\star}\beta_{1}\beta_{2}\frac{1}{k+1}\sum_{j=0}^{k}D_{j}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)}\beta_{1}.$
In the last line, we used $\frac{k+1}{N_{k}}\leq\beta_{1}$, $\frac{n_{j}}{N_{k}}=\frac{n_{j}}{k+1}\frac{k+1}{N_{k}}\leq\frac{n_{j}}{k+1}\beta_{1}\leq\beta_{1}\frac{\beta_{2}}{k+1}$ and $\gamma_{j}\leq\gamma^{\star}$ for all $k\geq 0,j\leq k$. Now we use the recursion assumption.
$D_{k+1}$ $\leq C_{\alpha}\frac{\lambda}{1-\lambda}\gamma^{\star}\beta_{1}\beta_{2}\frac{1}{k+1}\sum_{j=0}^{k}\frac{1}{(j+1)^{\alpha}}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)}\beta_{1}.$
Yet, $\sum_{j=0}^{k}\frac{1}{(j+1)^{\alpha}}\leq 1+B(\alpha)(k+1)^{1-\alpha}\leq 2B(\alpha)(k+1)^{1-\alpha}$, thus
$D_{k+1}$ $\leq 2C_{\alpha}\frac{\lambda}{1-\lambda}\gamma^{\star}\beta_{1}\beta_{2}B(\alpha)\frac{1}{(k+1)^{\alpha}}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)}\beta_{1}.$
$\leq 2C_{\alpha}\frac{\lambda}{1-\lambda}\gamma^{\star}\beta_{1}\beta_{2}B(\alpha)\frac{1}{(k+1)^{\alpha}}+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)^{\alpha}}\beta_{1}.$
Now, use that for all $k,(\frac{k+2}{k+1})^{\alpha}\leq 2^{\alpha}$:
$D_{k+1}$ $\leq 2^{1+\alpha}\frac{\beta_{1}}{(k+2)^{\alpha}}\left(C_{\alpha}\frac{\lambda}{1-\lambda}\gamma^{\star}\beta_{2}B(\alpha)+\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda}\right).$
Now using Eq. 10 gives:
$D_{k+1}$ $\leq\frac{\lambda}{1-\lambda}2^{2+\alpha}\gamma^{\star}\beta_{1}\beta_{2}B(\alpha)\frac{C_{\alpha}}{(k+2)^{\alpha}}.$
Now notice that $\lambda$ is such that
$\frac{\lambda}{1-\lambda}2^{2+\alpha}\gamma^{\star}\beta_{1}\beta_{2}B(\alpha)=1.$
This ensures that $D_{k+1}\leq\frac{C_{\alpha}}{(k+2)^{\alpha}}$ and concludes the recursion. ∎
Appendix F Experimental Design and Additional Results
F.1 Detailed Setup
General Setup.
Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work [15, 47]. Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies [41, 3], these policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost.
In this procedure, we need three splits: $D_{l}$ (of size $n_{l}$) to train the logging policy $\pi_{0}$, another split $D_{c}$ (of size $n_{c}$) to generate the logging feedback with $\pi_{0}$, and finally a test split $D_{test}$ (of size $n_{test}$) to compute the true risk $R(\pi)$ of any policy $\pi$. In our experiments, we split the training split $D_{train}$ (of size $N$) of the four datasets considered into $D_{l}$ ($n_{l}=0.05N$) and $D_{c}$ ($n_{c}=0.95N$) and use their test split $D_{test}$. The detailed statistics of the different splits can be found in Table 2. Recall that $K$ is the number of actions and $p$ the number of features.
All datasets are used without processing aside from CIFAR100 that we embed using a pretrained Resnet, obtaining a low dimensional vector of $100$ for each image.
Table 2: Detailed statistics of the datasets used.
Datasets $N$ $n_{l}$ $n_{c}$ $n_{test}$ $K$ $p$
MNIST 60 000 3000 57 000 10 000 10 784
FashionMNIST 60 000 3000 57 000 10 000 10 784
EMNIST 112 800 5640 107 160 18 800 47 784
CIFAR100 60 000 3000 57 000 10 000 100 500
Linear Gaussian Policies
We use the Linear Gaussian Policy of [41]. To obtain these policies, we restrict $f_{\theta}$ to:
$\forall x\in\mathcal{X},\quad f_{\theta}(x)=\operatorname*{arg\,max}_{a^{\prime}\in\mathcal{A}}\left\{x^{t}\theta_{a^{\prime}}\right\}$ (11)
This results in a parameter $\theta$ of dimension $d=p\times K$ with $p$ the dimension of the features $\phi(x)$ and $K$ the number of actions. We also restrict the family of distributions $\mathcal{Q}_{d+1}=\{Q_{\boldsymbol{\mu},\sigma}=\mathcal{N}(\boldsymbol{\mu},\sigma^{2}I_{d}),\boldsymbol{\mu}\in\mathbb{R}^{d},\sigma>0\}$ to independent Gaussians with shared scale. Estimating the propensity of $a$ given $x$ reduces the computation to a one dimensional integral:
$\pi_{\boldsymbol{\mu},\sigma}(a|x)$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1)}\left[\prod_{a^{\prime}\neq a}\Phi\left(\epsilon+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\right]$
with $\Phi$ the cumulative distribution function of the standard normal.
The starting behavior policy $\pi_{0}$.
$\pi_{0}$ is trained on $D_{l}$ (supervised manner) with the following parameters: We use $L_{2}$ regularization of $10^{-4}$. This is used to prevent the logging policy $\pi_{0}$ from being close to deterministic, allowing efficient learning with importance sampling. We use Adam [29] with a learning rate of $10^{-1}$ for $10$ epochs. Once it is trained, we use an inverse temperature parameter $\alpha$ on its score to interpolate between a uniform policy $\alpha=0$ and a trained policy $\alpha=1$.
Optimizing our learning objectives.
In each optimization subroutine, we use Adam [29] with a learning rate of $10^{-3}$ for $10$ epochs. The gradient of LIG policies is a one dimensional integral, and is approximated using $S=32$ samples.
$\pi_{\boldsymbol{\mu},\sigma}(a|x)$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1)}\left[\prod_{a^{\prime}\neq a}\Phi\left(\epsilon+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\right]$
$\approx\frac{1}{S}\sum_{s=1}^{S}\prod_{a^{\prime}\neq a}\Phi\left(\epsilon_{s}+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\quad\epsilon_{1},...,\epsilon_{S}\sim\mathcal{N}(0,1).$
F.2 More Experiments
Effect of the quality of $\pi_{0}$.
In the main experiments, we used $\alpha=0.2$, which makes the starting behavior policy keep some information about the good actions while being diffuse enough to put mass on all actions. We give here a more comprehensive view of how the algorithms behave depending on $\alpha$. The main take away, which reflects our first intuition, is that if we start far from $\pi_{\star}$, the improvement brought by adaptive off-policy approaches is substantial compared to using the static behavior policy $\pi_{0}$ to collect data.
Effect of Multiple Deployments while varying $\alpha$. We focus on CIFAR100 of large action space ($|\mathcal{A}|=100$). $\alpha=0.$ is the purely uniform behavior policy $\pi_{0}$, $\alpha=0.5$ is a very good policy but is still diffuse and $\alpha=1.$ is a peaked policy. Results are shown in Table 3. adaAdjLS outperforms adaLS in all scenarios, with a bigger gap when $\pi_{0}$ is far from a good policy.
Table 3: $R(\pi)$ varying $k$ and $\alpha$ in CIFAR100 (lower is better).
$\alpha$ $k$ adaLS (Alg 1) adaAdjLS (Alg 2)
$\alpha=0.$ 00 $-0.010$ $-0.010$
$1$ $-0.011$ $-0.011$
$5$ $-0.011$ $-0.011$
$10$ $-0.020$ $-0.026$
$100$ $-0.044$ $\mathbf{-0.077}$
$\alpha=0.5$ 00 $-0.268$ $-0.268$
$1$ $-0.430$ $-0.440$
$5$ $-0.500$ $-0.520$
$10$ $-0.527$ $-0.555$
$100$ $-0.583$ $\mathbf{-0.621}$
$\alpha=1.$ 00 $-0.531$ $-0.531$
$1$ $-0.569$ $-0.572$
$5$ $-0.590$ $-0.600$
$10$ $-0.600$ $-0.614$
$100$ $-0.630$ $\mathbf{-0.649}$
Effect of $\alpha$ on SCRM. We extend our evaluation of SCRM by varying the parameter $\alpha\in{0,0.5,1}$, with $k$ fixed at 10 across all datasets. When $\alpha=0$, the behavior policy is entirely uniform, making the learning task considerably more difficult—particularly in environments with large action spaces—since the policy provides minimal guidance and infrequently selects low-cost actions. Despite these challenges, both adaLS and adaAdjLS consistently outperform SCRM. As $\alpha$ increases, the behavior policy becomes more informative, improving SCRM’s performance. Nonetheless, our adaptive methods, especially adaAdjLS, continue to achieve superior results. The accompanying plots highlight the consistent advantage of adaAdjLS across all settings.
Figure 2: $R(\pi_{k})$ at each intermediate step, for $k=10$ and $\alpha=0.$.
Figure 3: $R(\pi_{k})$ at each intermediate step, for $k=10$ and $\alpha=0.5$.
Figure 4: $R(\pi_{k})$ at each intermediate step, for $k=10$ and $\alpha=1$.